# Are queries and keys always relevant? A case study on Transformer wave functions

### v1

@@ 4.1 Numerical experiments
where $\hat{\boldsymbol{S}}_{i}=(S_{i}^{x},S_{i}^{y},S_{i}^{z})$ and $J_{1},J_{2}\geq 0$ are antiferromagnetic couplings for nearest- and next-nearest neighbors, respectively. The ground state of this model exhibits magnetic order in the two distinct limits $J_{2}/J_{1}\ll 1$ and $J_{2}/J_{1}\gg 1$. Specifically, when $J_{2}=0$ ($J_{1}=0$) the model reduces to the unfrustrated Heisenberg model, characterized by long-range Néel (columnar) magnetic order [46, 47]. In the intermediate region, particularly around $J_{2}/J_{1}\approx 0.5$, the system becomes highly frustrated giving rise to exotic phases of matter [48]. The determination of the precise nature of the ground state in the frustrated region remains challenging and subject to debate [49, 50, 32].
We employ a Vision Transformer (ViT) wave function (see section 2.3) to approximate, in the VMC framework (see section 2.2), the ground state of this model on a $6\times 6$ lattice with periodic boundary conditions.
Table 1: Simulations on the $6\times 6$ lattice at $J_{2}/J_{1}=0.5$ with four layers ViT
@@
T5 -0.50318(1) 184,260 10h
Decoupled -0.50324(1) 184,260 10h
Factored -0.50321(1) 154,980 6h
In order to assess the efficacy of the three distinct attention mechanisms introduced in section 3.1, we perform simulations utilizing ViT architectures with identical hyperparameters (embedding dimension $d$, number of heads $h$, and number of layers $n_{l}$), modifying only the attention mechanism, namely T5 [see Eq. (4)], Decoupled [see Eq. (5)], and Factored [see Eq. (6)]. In Fig. 2, we report the optimization curves of the relative error of the variational energy with respect to the exact ground-state energy as a function of the optimization steps. On the left, we present the results for the unfrustrated case $(J_{2}/J_{1}=0)$ using ViT architectures with one [panel (a)] and four [panel (b)] layers. Instead, on the right we report the results in the frustrated regime ($J_{2}/J_{1}=0.5$), again using one [panel (c)] and four [panel (d)] layers architectures. We highlight that, while it is feasible to enhance the performance of the variational state by employing larger architectures (such as increasing the number of layers), utilizing T5 or Decoupled attention mechanisms with input-dependent attention weights, and subsequently increasing the computational complexity and parameter count via the matrices $Q$ and $K$, does not yield improved results compared to the Factored attention with input-independent attention weights. Notably, not only the final accuracies are practically identical, but also the learning dynamics exhibit similar behavior. In Table 1 we report the results of the ${J_{2}}$-${J_{1}}$ Heisenberg model on a $6\times 6$ lattice at ${J_{2}/J_{1}=0.5}$, obtained using a four-layer architecture (as in panel (c) of Fig. 2). The first column shows the final mean energy achieved by the different attention mechanisms, the second column indicates the number of parameters employed in the architectures, and the last column presents the computational time measured on a single GPU A100. The exact energy for this model can be computed using exact-diagonalization techniques, yielding $E_{0}/J_{1}=-0.503810$. It is worth noting that the accuracy of the results can be further enhanced by restoring the physical symmetries of the model through quantum number projection approaches [51, 52]; however, this goes beyond the scope of our work. The numerical simulations are performed in NetKet [53], with the VMC_SRt driver for optimizing the parameters. We used the implementation of the architecture based on the Vision Transformer from the UvA Deep Learning Tutorials [54], with the output layer described in section 2.3 and simple custom modifications for implementing the different attention mechanisms.
@@
Figure 3: a) Visualizations of the attention maps of a ViT with T5 attention mechanism [see Eq. (4)] for three different input spin configurations. When using initial random parameters there is a clear input dependence in the attention maps (top row). Instead, at the end of the optimization, the attention maps are practically input independent (bottom row). b) Visualizations of the input-dependent term (left panels) and of the input-independent term (right panels) of a ViT with Decoupled attention mechanism [see Eq. (5)]. After the optimization (bottom row), the input-dependent term is approximately the identity matrix shifted element-wise by a constant, thus Factored attention is recovered [see Eq. (6)]. In the plots, the input-dependent term has been averaged over $M=6000$ input configurations sampled from the optimized state. In both cases, the ViT architecture with a single layer $n_{l}=1$, embedding dimension $d=60$ and $h=10$ different heads is optimized on a $6\times 6$ lattice at $J_{2}/J_{1}=0.5$ (see panel (c) of Fig. 2). The plots are obtained by averaging the attention weights over all heads.
@@ 4.2 Analysis of the attention maps
The main result of the numerical simulations reported in Fig. 2 and discussed in section 4.1 is that, using a ViT employing T5, Decoupled and Factored attention, the final accuracy is practically the same (see Table 1). This suggests that, in the case of T5 and Decoupled attention, queries and keys are effectively not used in the optimized solution. To validate this statement we study the attention maps. For the analysis, we use a single-layer architecture, where the interpretation of the results is simplified since the patches are only mixed within the attention mechanism, and the subsequent MLP cannot modify the relative weights among the various attention vectors. In panel $a)$ of Fig. 3, we consider the case of T5 attention, plotting the attention weights defined in Eq. (4) for three different input spin configurations. We first check that at the beginning, with random parameters, the attention maps depend on the inputs (top row), ensuring that we have an unbiased initialization. In the bottom row, we show that the architecture after optimization produces input-independent attention maps, thus automatically recovering a positional only solution. In panel $b)$ of Fig. 3, we consider the case of Decoupled attention, plotting separately the input dependent and the positional contributions of the attention weights [see Eq. (5)]. Again, after optimization the network swaps from an unbiased solution (top row) to a positional only solution (bottom row), where the input-dependent term converges approximately to the identity matrix shifted element-wise by a constant. In other words, Factored attention is spontaneously recovered from the Decoupled version (see section 3.1). Interestingly, a related result, but in a completely different application, has been found in the recent work of Ref. [55].
@@ 4.3 Representation of physical ground states with Factored attention
In this section, we provide analytic calculations about the efficacy of input-independent attention mechanisms for approximating quantum states. We first examine an analytical solvable quantum many-body Hamiltonian, developing an exact mapping between its ground state and a single layer of two-headed Factored attention. Building upon this result, we extend our analysis to scenarios where the ground state lacks analytical solutions, providing insights into why attention mechanisms including queries and keys [as in Eq. (4) and Eq. (5)] should converge to positional only solutions when studying large systems. Interestingly, the same arguments can be applied to other domains, such as NLP or computer vision.
As illustrative example of solvable quantum many-body Hamiltonian, we consider the Shastry-Sutherland model [56], which captures the low-temperature properties of $\text{SrCu}_{2}(\text{BO}_{3})_{2}$, a compound known for its intriguing physical properties [57]. In a finite range of the frustration ratio, the ground state of this model is represented as a product of singlets between next-nearest-neighbor spins arranged on a square lattice [56], refer to Fig. 4 for a graphical representation. Here, we want to show that a single-layer ViT with Factored attention [see Eq. (6)] can represent exactly this ground state. Working on a $L\times L$ square lattice with periodic boundary conditions, we partition input spin configurations into $b\times b$ patches, with $b=2$ (see Fig. 4), which are then flattened to construct input sequences. Assuming an embedding dimension of $d=b^{2}=4$ and choosing the embedding matrix to be the identity, the $i$-th input vector is $\boldsymbol{x}_{i}=(\sigma_{i,1},\sigma_{i,2},\sigma_{i,3},\sigma_{i,4})^{T}$, where $i=1,\dots,L^{2}/b^{2}$.
@@
Then, we apply the Multi-Head attention mechanism [1] with $h=2$ heads. Considering the value matrices:
@@
Now, we assume the $L^{2}/b^{2}\times L^{2}/b^{2}$ attention matrices to be $\alpha^{(1)}_{ij}=\delta_{i,j}$ and $\alpha^{(2)}_{ij}=\delta_{i,S(i)}$, where:
$S(i)=\begin{cases}(i+1)\%(L^{2}/b^{2})&\text{if}\ \ i\%(L/b)=0,\\ (i+L/b)\%(L^{2}/b^{2})+1&\text{otherwise,}\end{cases}$ (10)
@@
The hidden representation is obtained by summing all the output vectors $\boldsymbol{z}=\sum_{i=1}^{L^{2}/b^{2}}\boldsymbol{y}_{i}$, where $\boldsymbol{z}\in\mathbb{R}^{d}$:
$\boldsymbol{z}=\left(\sum_{i=1}^{L^{2}/b^{2}}F(V_{11}^{(1)}\sigma_{i,4}+V_{14}^{(2)}\sigma_{S(i),1}+c),\sum_{i=1}^{L^{2}/b^{2}}F(V_{22}^{(1)}\sigma_{i,2}+V_{23}^{(1)}\sigma_{i,3}+c),0,0\right)^{T}.$ (15)
@@
$\text{Log}[\Psi_{\theta}(\sigma)]=\sum_{i=1}^{L^{2}/b^{2}}\left[F(V_{11}^{(1)}\sigma_{i,4}+V_{14}^{(2)}\sigma_{S(i),1}+c)+F(V_{22}^{(1)}\sigma_{i,2}+V_{23}^{(1)}\sigma_{i,3}+c)\right]\ .$ (16)
@@
$\lim_{|i-j|\to+\infty}\braket{\hat{A}_{i}\hat{A}_{j}}=\braket{\hat{A}_{i}}\braket{\hat{A}_{j}}\ ,$ (18)
where $\hat{A}_{i}$ is a generic local operator. According to the cluster property, correlations must decay with distance and, in the thermodynamic limit, sites that are infinitely distant become uncorrelated. As shown in the previous mapping, the role of the attention weights is to connect correlated inputs. Therefore, for systems for which the property in Eq. (18) holds, we expect the attention weights connecting spins far apart in the system to be close to zero, regardless the specific values of the spins. Interestingly, to reproduce this long-distance behavior using standard T5 [Eq. (4)] or Decoupled [see Eq. (5)] attention mechanisms we have to require $Q=K=0$. In other words, the standard attention mechanisms should converge to positional only solutions, thereby to the Factored version [see Eq. (6)]. This argument, which exploits only the correlations among the elements of the input sequence, can be extended to any domain provided that the input sequence is long enough that correlations decay significantly within the scale of the system. For example, even in NLP or in computer vision tasks, when considering long input sentences or large images, it must be true that words or patch of pixels that are really far apart are uncorrelated, and so in this limit queries and keys should be optimized to zero. However, when dealing with finite sequences, this argument can have a marginal impact, and using input dependent attention weights as in Eq. (4) can provide a good inductive bias for solving the task.
@@ 5 Conclusion
In this work we showed that, when training a Transformer to approximate ground states of quantum many-body Hamiltonians, the standard attention mechanism with relative positional encoding [40] yields equivalent performance of a simplified version, the Factored attention. The latter utilizes input-independent attention weights, resulting in fewer parameters and reduced computational cost. Moreover, starting from analytical computations, we established a direct link between attention weights and correlations. We observed that, if the dominant correlation lengths necessary to solve a specific task are shorter than the total input size, the weights in conventional attention mechanisms (e.g., T5 [40]) should converge towards input-independent solutions. Interestingly, the same considerations can be extended in NLP and computer vision domains. For instance, in image classification tasks, the pertinent scale is associated to the extension of the objects requiring detection, typically smaller than the entire image. A straightforward approach to mitigate the potential issues associated to the relationship between long-range behavior of the correlations and queries and keys is the implementation of local attention mechanisms, wherein attention weights beyond a specified distance are manually set to zero. In Ref. [61] has been found that it is possible to use short-range attention for the majority of layers in the Transformer and recover the same performance of long-range language modelling. An interesting future direction of research could be the design of attention mechanisms that are able to describe the decay of long-range correlations without the necessity to set queries and keys to zero or without employing local attention mechanisms.

### v2

@@ 4.1 Numerical experiments
where $\hat{\boldsymbol{S}}_{i}=(S_{i}^{x},S_{i}^{y},S_{i}^{z})$ and $J_{1},J_{2}\geq 0$ are antiferromagnetic couplings for nearest- and next-nearest neighbors, respectively. The ground state of this model exhibits magnetic order in the two distinct limits $J_{2}/J_{1}\ll 1$ and $J_{2}/J_{1}\gg 1$. Specifically, when $J_{2}=0$ ($J_{1}=0$) the model reduces to the unfrustrated Heisenberg model, characterized by long-range Néel (columnar) magnetic order [47, 48]. In the intermediate region, particularly around $J_{2}/J_{1}\approx 0.5$, the system becomes highly frustrated, giving rise to exotic phases of matter [49]. The determination of the precise nature of the ground state in the frustrated region remains challenging and subject to debate [50, 51, 33].
We employ a ViT wave function (see section 2.3) to approximate, in the VMC framework (see section 2.2), the ground state of this model on a $L\times L$ lattice with periodic boundary conditions. In order to assess the efficacy of the three distinct attention mechanisms introduced in section 3.1, we perform simulations on a $6\times 6$ cluster utilizing ViT architectures with identical hyperparameters (embedding dimension $d$, number of heads $h$, number of layers $n_{l}$, and linear patch size $b$), modifying only the attention mechanism, namely T5 [see Eq. (4)], Decoupled [see Eq. (5)], and Factored [see Eq. (6)]. In Fig. 2, we report the optimization curves of the relative error of the variational energy with respect to the exact ground-state energy as a function of the optimization steps. On the left, we present the results for the unfrustrated case $(J_{2}/J_{1}=0)$ using ViT architectures with one [panel (a)] and four [panel (b)] layers. Instead, on the right, we report the results in the frustrated regime ($J_{2}/J_{1}=0.5$), again using one [panel (c)] and four [panel (d)] layers architectures. We emphasize that, although it is possible to enhance the performance of the variational state by employing larger architectures, such as increasing the number of layers, considering larger embedding dimensions or augmenting the number of heads [15, 16, 19], the use of T5 or Decoupled attention mechanisms with input-dependent attention weights, and the subsequent increase of computational complexity and parameter count via the matrices $Q$ and $K$, does not produce improved results compared to Factored attention with input-independent attention weights. Notably, not only are the final accuracies practically identical, but also the learning dynamics exhibit similar behavior.
Figure 3: Left panel: Relative error $\Delta\varepsilon=|(E_{0}-E_{\text{ViT}})/E_{0}|$ at $J_{2}/J_{1}=0.5$ as a function of the system size for ViT architectures with the three different attention mechanisms, namely Factored (orange circles), Decoupled (green squares) and T5 (blue diamonds). The reference ground state energies are taken from exact diagonalization for $L=6$ $(-0.503810)$ [52] and from variance extrapolation for $L=8$ ($-0.49906$[50] and $L=10$ ($-0.497715$[30]. Right panel: Time per optimization step (in seconds) measured on a single GPU A100 for the three attention mechanisms as a function of the system size. For all simulations a ViT architecture with hyperparameters $d=10$, $h=10$, $b=2$ and $n_{l}=4$ is considered. The model is optimized using the SR optimization method for $5\times 10^{3}$ steps, employing $M=6\times 10^{3}$ training samples (see section 2.2). A cosine decay learning rate scheduler is applied, starting with an initial value of $\tau=0.03$.
In Fig. 3, we extend our analysis to larger system sizes, specifically for $L=8$ and $L=10$. We focus on an architecture with the same hyperparameters for the different sizes: number of heads ($h=10$), embedding dimension ($d=60$), linear patch size ($b=2$) and number of layers ($n_{l}=4$). The left panel displays the relative error of the variational energy as a function of the system size $L$ at $J_{2}/J_{1}=0.5$. The reference energies used to compute the accuracy are obtained through exact diagonalization for $L=6$ [52] and through variance extrapolation from Ref. [50] and Ref. [30], for $L=8$ and $L=10$, respectively. This plot demonstrates that the accuracy remains size-consistent across the tested clusters, showing a constant behavior when increasing the system size, despite the fact that the network has fixed complexity. In the right panel, we present the computational time per optimization step as a function of the system size measured on a single GPU A100. The data illustrate how the efficiency gap between the Factored attention mechanism and the other attention mechanisms becomes more pronounced when increasing the system size.
@@
T5 -0.503182(9) 184,260 10h
Decoupled -0.503243(9) 184,260 10h
Factored -0.503216(8) 154,980 6h
Energy Parameters Time
T5 -0.497025(6) 184,900 28h
Decoupled -0.497108(6) 184,900 28h
Factored -0.497184(6) 155,620 12.5h
Table 1: Results for the ${J_{2}}$-${J_{1}}$ Heisenberg model at $J_{2}/J_{1}=0.5$ obtained using a ViT architecture with a number of heads $h=10$, embedding dimension $d=60$, linear patch size $b=2$ and a number of layers $n_{l}=4$ on a $6\times 6$ (left) and on a $10\times 10$ lattice (right).
In Table 1 we report the results on a $6\times 6$ and a $10\times 10$ lattice at ${J_{2}/J_{1}=0.5}$, obtained using a four-layer architecture. In both tables, the first column shows the final mean energy achieved by the different attention mechanisms, the second column indicates the number of parameters employed in the architectures, and the last column presents the total computational time measured on a single GPU A100 to perform $5\times 10^{3}$ optimization steps. It is worth noting that the accuracy of the results can be further enhanced by restoring the physical symmetries of the model through quantum number projection approaches [53, 54]; however, this goes beyond the scope of our work.
@@
Figure 4: Panel a: Visualizations of the attention maps of a ViT with T5 attention mechanism [see Eq. (4)] for three different input spin configurations. When using initial random parameters there is a clear input dependence in the attention maps (top row). Instead, at the end of the optimization, the attention maps are practically input independent (bottom row). Panel b: Visualizations of the input-dependent term (left panels) and of the input-independent term (right panels) of a ViT with Decoupled attention mechanism [see Eq. (5)]. After the optimization (bottom row), the input-dependent term is approximately the identity matrix shifted element-wise by a constant, thus Factored attention is recovered [see Eq. (6)]. In the plots, the input-dependent term has been averaged over $M=6\times 10^{3}$ input configurations sampled from the optimized state. The presented results are obtained by optimizing a ViT architecture with a single layer $n_{l}=1$, embedding dimension $d=60$ and $h=10$ different heads on a $6\times 6$ lattice at $J_{2}/J_{1}=0.5$ (see panel (c) of Fig. 2). The linear patch size is taken to be $b=2$, thus we have $n=9$ patches and the resulting attention maps have shape $9\times 9$. The plots are obtained by averaging the attention weights over all heads.
@@ 4.2 Analysis of the attention maps
The main result of the numerical simulations reported in Figs. 23 and discussed in section 4.1 is that, using a ViT employing T5, Decoupled and Factored attention, the final accuracy is practically the same (see Table 1). This suggests that, in the case of T5 and Decoupled attention, queries and keys are effectively not used in the optimized solution. To validate this statement, we study the attention maps. For the analysis, we used a single-layer architecture, where the interpretation of the results is simplified since the patches are only mixed within the attention mechanism, and the subsequent MLP cannot modify the relative weights among the various attention vectors. In panel $(a)$ of Fig. 4, we consider the case of T5 attention, plotting the attention weights defined in Eq. (4) for three different input spin configurations. We first check that at the beginning, with random parameters, the attention maps depend on the inputs (top row), ensuring that we have an unbiased initialization. In the bottom row, we show that the architecture after optimization produces input-independent attention maps, thus automatically recovering a positional-only solution. In panel $(b)$ of Fig. 4, we consider the case of Decoupled attention, plotting separately the input dependent and the positional contributions of the attention weights [see Eq. (5)]. Again, after optimization the network swaps from an unbiased solution (top row) to a positional only solution (bottom row), where the input-dependent term converges approximately to the identity matrix shifted element-wise by a constant. In other words, Factored attention is spontaneously recovered from the Decoupled version (see section 3.1).
@@ 4.3 Representation of physical ground states with Factored attention
In this section, we provide analytic calculations about the efficacy of input-independent attention mechanisms for approximating quantum states. We first examine an analytically solvable quantum many-body Hamiltonian, developing an exact mapping between its ground state and a single layer of two-headed Factored attention. Building upon this result, we extend our analysis to scenarios where the ground state lacks analytical solutions, providing insights into why attention mechanisms including queries and keys [as in Eq. (4) and Eq. (5)] should converge to positional-only solutions when studying large systems.
@@
As an illustrative example of a solvable quantum many-body Hamiltonian, we consider the Shastry-Sutherland model [55], which captures the low-temperature properties of $\text{SrCu}_{2}(\text{BO}_{3})_{2}$, a compound known for its intriguing physical properties [56]. In a finite range of the frustration ratio, the ground state of this model is represented as a product of singlets between next-nearest-neighbor spins arranged on a square lattice [55], refer to Fig. 5 for a graphical representation. Here, we want to show that a single-layer ViT with Factored attention [see Eq. (6)] can represent exactly this ground state. Working on a $L\times L$ square lattice with periodic boundary conditions, we partition input spin configurations into $b\times b$ patches, with $b=2$ (see Fig. 5), which are then flattened to construct input sequences. Assuming an embedding dimension of $d=b^{2}=4$ and choosing the embedding matrix to be the identity, the $i$-th input vector is $\boldsymbol{x}_{i}=(\sigma_{i,1},\sigma_{i,2},\sigma_{i,3},\sigma_{i,4})^{T}$, where $i=1,\dots,n$, with $n=L^{2}/b^{2}$. Then, we apply the Multi-Head attention mechanism [1] with $h=2$ heads. Considering the value matrices:
@@
Now, we assume the $n\times n$ attention matrices to be $\alpha^{(1)}_{ij}=\delta_{i,j}$ and $\alpha^{(2)}_{ij}=\delta_{i,S(i)}$, where:
$S(i)=\begin{cases}(i+1)\%n&\text{if}\ \ i\%(L/b)=0,\\ (i+L/b)\%n+1&\text{otherwise,}\end{cases}$ (10)
@@
The hidden representation is obtained by summing all the output vectors $\boldsymbol{z}=\sum_{i=1}^{n}\boldsymbol{y}_{i}$, where $\boldsymbol{z}\in\mathbb{R}^{d}$:
$\boldsymbol{z}=\left(\sum_{i=1}^{n}F(V_{11}^{(1)}\sigma_{i,4}+V_{14}^{(2)}\sigma_{S(i),1}+c),\sum_{i=1}^{n}F(V_{22}^{(1)}\sigma_{i,2}+V_{23}^{(1)}\sigma_{i,3}+c),0,0\right)^{T}.$ (15)
@@
$\text{Log}[\Psi_{\theta}(\sigma)]=\sum_{i=1}^{n}\left[F(V_{11}^{(1)}\sigma_{i,4}+V_{14}^{(2)}\sigma_{S(i),1}+c)+F(V_{22}^{(1)}\sigma_{i,2}+V_{23}^{(1)}\sigma_{i,3}+c)\right]\ .$ (16)
@@
$\lim_{|i-j|\to+\infty}\braket{\hat{B}_{i}\hat{B}_{j}}=\braket{\hat{B}_{i}}\braket{\hat{B}_{j}}\ ,$ (18)
where $\hat{B}_{i}$ is a generic local operator. According to the cluster property, correlations must decay with distance and, in the thermodynamic limit, sites that are infinitely distant become uncorrelated. As shown in the previous mapping, the role of the attention weights is to connect correlated inputs. Therefore, for systems for which the property in Eq. (18) holds, we expect the attention weights connecting spins far apart in the system to be close to zero, regardless of the specific values of the spins. Interestingly, to reproduce this long-distance behavior using standard T5 [Eq. (4)] or Decoupled [see Eq. (5)] attention mechanisms we have to require $Q=K=0$. In other words, the standard attention mechanisms should converge to positional only solutions, thereby to the Factored version [see Eq. (6)]. This argument, which exploits only the correlations among the elements of the input sequence, can be extended to any domain provided that the input sequence is long enough that correlations decay significantly within the scale of the system. For example, even in NLP or in computer vision tasks, when considering long input sentences or large images, it must be true that words or patches of pixels that are really far apart are uncorrelated, and so in this limit queries and keys should be optimized to zero. However, when dealing with finite sequences, this argument can have a marginal impact, and using input dependent attention weights as in Eq. (4) can provide a good inductive bias for solving the task.
@@ 5 Conclusion
In this work, we showed that, when training a Transformer to approximate ground states of quantum many-body Hamiltonians, the standard attention mechanism yields equivalent performance to a simplified version, the Factored attention. The latter utilizes input-independent attention weights, resulting in fewer parameters and reduced computational cost. Moreover, starting from analytical computations, we established a direct link between attention weights and correlations. We observed that if the dominant correlation lengths necessary to solve a specific task are shorter than the total input size, the weights in conventional attention mechanisms (e.g., T5 [41]) should converge towards input-independent solutions. Interestingly, the same considerations can be extended to NLP and computer vision domains. For example, in image classification tasks, the pertinent scale is associated with the extension of objects requiring detection, typically smaller than the entire image. A straightforward approach to mitigate potential problems associated with the relationship between long-range behavior of correlations and queries and keys is the implementation of local attention mechanisms, wherein attention weights beyond a specified distance are manually set to zero. In Ref. [60] it has been found that it is possible to use short-range attention for the majority of layers in the Transformer and recover the same performance of long-range language modeling. However, we emphasize that a necessary condition for the validity of our results is the possibility to partition the input sequences into effectively uncorrelated segments. This requirement may not hold universally across NLP applications. For instance, studies have demonstrated that correlations can extend over arbitrarily long scales in literary texts [61, 62], and that, for specific tasks, global token mechanisms are preferred [63]. An interesting future direction of research could be the design of attention mechanisms that are able to describe the decay of long-range correlations without the necessity to set queries and keys to zero or without employing local attention mechanisms.
Reproducibility
The variational quantum Monte Carlo and the ViT architecture were implemented in JAX [24]. The implementation of the Stochastic Reconfiguration [15] is available on NetKet [64] under the name of VMC_SRt. The ViT architecture used in this paper is available at https://zenodo.org/records/14060431.